# Track-On2: Enhancing Online Point Tracking with Memory

### v1

@@ 4.4 Efficiency
We report inference speed (frames per second, FPS) and maximum GPU memory usage as a function of inference-time memory size $L_{i}$ for tracking 64, 256, and 1024 points in Fig. 12. All results are obtained with uniform grid queries on the first frame, using FP32 on a single NVIDIA A100 GPU. Our model can track 64 sparse points at over 35 FPS while using only 0.64 GB of GPU memory. With 256 points, it sustains nearly 35 FPS with less than 1 GB memory, even as $L_{i}$ increases. For 1024 points, the maximum memory usage remains only 1.5 GB at $L_{i}=72$, while keeping FPS above 20. Since state-of-the-art accuracy on longer videos is achieved with $L_{i}=72$, these results demonstrate that our model satisfies real-time criteria at more than 30 FPS in practical configurations.
@@ 4.5 Ablation Study
    choice of DINO backbone variants (Sec. III);
@@
    contributions of model components (Sec. IV);
@@ 4.5.1 DINO as Backbone
TABLE III: DINO backbones in ViT-Adapter. We replace the DINOv2 backbone in our ViT-Adapter with DINOv3. All metrics are higher-is-better.
@@
DINOv3 was introduced as a scaled-up version of DINOv2, with larger model capacity and training data, and has shown strong gains across many vision tasks. To study its effect in our framework, we evaluate different backbone options in Table III. Specifically, we include DINOv2 ViT-S, DINOv3 ViT-S, and DINOv3 ViT-S+. We restrict our study to these smaller variants to maintain efficiency. Since DINOv2 uses a patch size of 14 and DINOv3 uses 16, we upsample the ViT inputs (but not the ViT-Adapter) with a ratio of $16/14$ to match the token count, and apply this upsampling consistently in our main model. Interestingly, despite being trained on a much larger dataset, DINOv3 does not outperform DINOv2 when using the same ViT-S architecture: DINOv2 achieves 0.6 and 0.2 higher AJ scores on DAVIS and RoboTAP, respectively. We attribute this to model capacity, as DINOv3 was primarily designed for large-scale backbones such as ViT-7B, while smaller variants are distilled from it. However, moving from ViT-S (21M parameters) to ViT-S+ (29M, with SwiGLU feedforward layers instead of standard MLPs) yields consistent gains, surpassing DINOv2 on RoboTAP while remaining on par on DAVIS.
@@ 4.5.2 Components
TABLE IV: Model Components. Removing individual components of our model, namely memory ($\mathbf{M}$) the re-ranking module ($\Phi_{\text{re-rank}}$), and offset head ($\Phi_{\text{off}}$), one at a time. All metrics are higher-is-better.
@@
We conducted an experiment to examine the impact of each main component, by removing them one at a time while keeping other modules unchanged. First, we removed the memory $\mathbf{M}$. Second, we removed the re-ranking module $\Phi_{\text{re-rank}}$. Lastly, we removed the offset head $\Phi_{\text{off}}$, eliminating the calculation of additional offsets. Instead, we used the coarse prediction, i.e. the selected patch center, as the final prediction. Note that, we do not apply inference-time memory extension to models in this comparison. From the results in Table IV, we can make the following observations: (i) Memory is critical; without it, the model operates only at the frame level and loses all temporal context, leading to AJ drops of 10.0 and 4.7 points on DAVIS and RoboTAP, respectively, and reducing OA on DAVIS from 91.5 to 81.5. (ii) The re-ranking module improves all metrics, notably increasing AJ by 4.7 on DAVIS and 2.7 on RoboTAP, as it introduces specialized queries that refine correspondences. (iii) The offset head is crucial for fine-grained localization, with $\delta^{x}_{avg}$dropping by $\sim$6 points on DAVIS and 5.6 on RoboTAP when removed. Without offsets, the model defaults to predicting patch centers; however, unlike memory or re-ranking, this does not directly affect the quality of the queries, but instead controls the granularity of localization. Interestingly, OA slightly improves in this case, likely because removing the offset loss term in Eq. (7) rebalances optimization, placing more emphasis on occlusion prediction.
@@ 4.5.3 Video Length and Memory Size
TABLE V: Effect of training video length and memory size. We compare different training video lengths $T$ and memory sizes $L$, without applying inference-time memory extension (i.e. $L=L_{i}$). Longer training videos consistently improve performance across datasets and are especially critical for robustness in very long PO sequences. All metrics are higher-is-better.
Model $T$ $L$ DAVIS RoboTAP PO
@@
A 24 18 77.5 90.2 76.0 89.0 46.8
B 36 18 78.3 90.7 78.2 91.2 51.3
C 48 18 79.0 90.7 79.1 92.2 51.5
D 48 24 79.1 91.5 78.9 91.5 51.5
Results in Table V show that training video length is crucial. Increasing $T$ while keeping memory size fixed (models A $\rightarrow$ B $\rightarrow$ C, with $T=24\rightarrow 36\rightarrow 48$ and $L=18$) consistently improves performance across all datasets and metrics, with especially large gains in PO Survival. For example, OA on RoboTAP improves by 3.2 points when increasing $T$ from 24 to 48. By contrast, increasing memory size from $L=18\rightarrow 24$ (model D) yields modest improvements in DAVIS but slightly reduces performance in RoboTAP, while PO remains unchanged, suggesting that memory capacity alone is less critical than providing sufficiently long training sequences. Finally, we note that the full potential of memory is unlocked when it is extended at inference time, which we discuss in the next section.
@@ 4.5.5 Training Video Frame Sampling
While we sample $T$ frames from the 120-frame videos in the dataset, the way these frames are selected can be important. For example, if the object or camera motion is slow, sampling only sequential frames (stride 1) may provide little useful signal for learning motion. To explore the optimal frame sampling strategy, we experiment with two sampling methods and different ratios, as shown in Table VI.
@@
TABLE VI: Training video frame sampling. Comparison of uniform and random strategies with different ratios $r$, where $r$ controls the temporal crop size before selecting $T$ training frames. All metrics are higher-is-better.
@@ 5 Conclusion
Limitations and Future Work: Despite its strengths, our model has several limitations. Track-On2 prioritizes appearance over motion, as patch classification relies entirely on feature similarity, and memory captures motion only in an implicit manner. This can lead to visually unstable predictions, such as slight jittering in regions with uniform texture, where consistent patch localization becomes more ambiguous. To address this, explicitly modeling motion within the architecture is a promising future direction. Such a design would increase the model’s motion sensitivity while reducing over-reliance on appearance. In addition, future work includes exploring real-world fine-tuning and improving pseudo-labeling strategies to further enhance robustness and generalization.
@@ 6 Acknowledgements
This project is funded by the European Union (ERC, ENSURE, 101116486) with additional compute support from Leonardo Booster (EuroHPC Joint Undertaking, EHPC-AI-2024A01-060). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Weidi Xie would like to acknowledge the Scientific Research Innovation Capability Support Project for Young Faculty (ZYGXQNJSKYCXNLZCXM-I22).

### v2

@@ 4.4 Efficiency
We report inference speed (frames per second, FPS) and peak GPU memory usage as a function of the inference-time memory size $L_{i}$ for tracking 64, 256, and 1024 points in Fig. 14. All results are obtained using uniform grid queries initialized on the first frame, evaluated in FP32 on a single NVIDIA A100 GPU.
For sparse tracking with 64 points, our model operates at over 35 FPS while consuming only 0.37 GB of GPU memory. When tracking 256 points, it maintains close to 35 FPS with less than 0.5 GB of memory, even as $L_{i}$ increases. For denser tracking with 1024 points, the peak memory usage remains around 1.2 GB at $L_{i}{=}72$, while sustaining approximately 20 FPS. Since state-of-the-art accuracy on longer videos is achieved with higher memory sizes, these results demonstrate that our model meets real-time requirements for practical sparse-point tracking scenarios while maintaining a low memory footprint.
@@ 4.5 Ablation Study
    architectural and training differences between Track-On and Track-On2 (Sec. 4.5.1).
@@
    choice of DINO backbone variants (Sec. IV);
@@
4.5.1 Track-On to Track-On2
TABLE III: Ablation from Track-On to Track-On2. Dv3 indicates a DINOv3 backbone; FPN denotes multi-scale feature fusion; Mem. denotes a unified single-memory design; Data indicates training with the Track-On2 data and sampling strategy; $T$ and $L$ are the training clip length and training-time memory size. Row (i), shaded in gray, corresponds to the original Track-On trained with the original data regime; all other variants use the Track-On2 training setup.
  Dv3 FPN Mem. Data $T$ $L$ DAVIS RoboTAP PO
            $\delta^{x}_{avg}$  OA    $\delta^{x}_{avg}$  OA    Survival
(i)    ✗    ✗    ✗    ✗    24    12 78.0    90.8    76.4    89.4    49.5
(ii)    ✗    ✗    ✗    ✓    24    12 78.9    91.3    79.4    92.8    53.9
(iii)    ✗    ✗    ✗    ✓    48    12 79.1    91.5    80.0    92.9    57.5
(iv)    ✗    ✗    ✗    ✓    48    24 Training OOM
(v)    ✗    ✗    ✓    ✓    24    12 78.8    90.5    79.3    92.1    50.8
(vi)    ✗    ✗    ✓    ✓    48    12 78.9    91.5    80.6    92.8    54.4
(vii)    ✗    ✗    ✓    ✓    48    24 78.7    91.8    79.9    92.8    57.4
(viii)    ✗    ✓    ✓    ✓    48    24 79.8    91.7    80.0    92.7    57.8
(ix)    ✓    ✓    ✓    ✓    48    24 79.9    92.0    80.5    93.4    57.7
We conduct a controlled ablation study in Table III to analyze the design changes from Track-On to Track-On2. Starting from the original Track-On architecture, we progressively modify one component at a time while keeping all other settings fixed.
We first isolate the effect of the new training regime alone. Comparing the original Track-On baseline with the same architecture trained under the Track-On2 training setup, using the updated dataset and revised frame sampling strategy, shows consistent improvements across datasets, indicating that a substantial portion of the gains arises from improved training rather than architectural modifications.
We next examine the impact of memory design under different training horizons. With short training clips (T=24T{=}24), the dual-memory formulation used in Track-On performs better than a unified-memory design (Sec. 3.2.1), consistent with prior observations. However, when the training horizon is extended to T=48T{=}48, this performance gap largely disappears, suggesting that longer-context training reduces the need for an additional spatial memory. Importantly, increasing the training-time memory size to L=24L{=}24 renders the dual-memory formulation infeasible due to out-of-memory errors on 64GB GPUs, whereas the unified-memory design remains stable and scalable. This enables more efficient training and inference, reducing GPU memory usage by approximately $20\%$ and improving inference speed by over 5 FPS when tracking 256 points with $L_{i}{=}72$.
Finally, we evaluate the remaining architectural refinements. Introducing FPN-style multi-scale feature fusion (Sec. 3.2.2) and replacing DINOv2 with DINOv3 yields the final Track-On2 model with the strongest overall performance.
Overall, these results show that the primary gains in Track-On2 arise from improved training with longer temporal context, while the architectural choices, particularly unified memory and multi-scale feature fusion, are crucial for making such training feasible, efficient, and scalable. While accuracy gains over Track-On may be modest on short sequences, Track-On2 consistently achieves comparable or improved robustness with substantially better efficiency, especially on long videos.
@@ 4.5.2 DINO as Backbone
TABLE IV: DINO backbones in ViT-Adapter. We replace the DINOv2 backbone in our ViT-Adapter with DINOv3. All metrics are higher-is-better.
@@
DINOv3 was introduced as a scaled-up version of DINOv2, with larger model capacity and training data, and has shown strong gains across many vision tasks. To study its effect in our framework, we evaluate different backbone options in Table IV. Specifically, we include DINOv2 ViT-S, DINOv3 ViT-S, and DINOv3 ViT-S+. We restrict our study to these smaller variants to maintain efficiency. Since DINOv2 uses a patch size of 14 and DINOv3 uses 16, we upsample the ViT inputs (but not the ViT-Adapter) with a ratio of $16/14$ to match the token count, and apply this upsampling consistently in our main model. Interestingly, despite being trained on a much larger dataset, DINOv3 does not outperform DINOv2 when using the same ViT-S architecture: DINOv2 achieves 0.6 and 0.2 higher AJ scores on DAVIS and RoboTAP, respectively. We attribute this to model capacity, as DINOv3 was primarily designed for large-scale backbones such as ViT-7B, while smaller variants are distilled from it. However, moving from ViT-S (21M parameters) to ViT-S+ (29M, with SwiGLU feedforward layers instead of standard MLPs) yields consistent gains, surpassing DINOv2 on RoboTAP while remaining on par on DAVIS.
@@ 4.5.3 Components
TABLE V: Model Components. Removing individual components of our model, namely memory ($\mathbf{M}$) the re-ranking module ($\Phi_{\text{re-rank}}$), and offset head ($\Phi_{\text{off}}$), one at a time. All metrics are higher-is-better.
@@
We conducted an experiment to examine the impact of each main component, by removing them one at a time while keeping other modules unchanged. First, we removed the memory $\mathbf{M}$. Second, we removed the re-ranking module $\Phi_{\text{re-rank}}$. Lastly, we removed the offset head $\Phi_{\text{off}}$, eliminating the calculation of additional offsets. Instead, we used the coarse prediction, i.e. the selected patch center, as the final prediction. Note that, we do not apply inference-time memory extension to models in this comparison. From the results in Table V, we can make the following observations: (i) Memory is critical; without it, the model operates only at the frame level and loses all temporal context, leading to AJ drops of 10.0 and 4.7 points on DAVIS and RoboTAP, respectively, and reducing OA on DAVIS from 91.5 to 81.5. (ii) The re-ranking module improves all metrics, notably increasing AJ by 4.7 on DAVIS and 2.7 on RoboTAP, as it introduces specialized queries that refine correspondences. (iii) The offset head is crucial for fine-grained localization, with $\delta^{x}_{avg}$dropping by $\sim$6 points on DAVIS and 5.6 on RoboTAP when removed. Without offsets, the model defaults to predicting patch centers; however, unlike memory or re-ranking, this does not directly affect the quality of the queries, but controls the granularity of localization. Interestingly, OA slightly improves in this case, likely because removing the offset loss term in Eq. (7) rebalances optimization, placing more emphasis on occlusion prediction.
@@ 4.5.4 Video Length and Memory Size
Results in Table VI show that training clip length is the dominant factor, while memory capacity provide complementary gains. Increasing $T$ while keeping memory size fixed (models i $\rightarrow$ $i$ $\rightarrow$ iii, with $T=24\rightarrow 36\rightarrow 48$ and $L=18$) consistently improves performance across all datasets and metrics, with especially large gains in PO Survival. For example, OA on RoboTAP improves by 3.2 points when increasing $T$ from 24 to 48. By contrast, increasing memory size from $L=18\rightarrow 24$ (model iv) yields modest improvements in DAVIS but slightly reduces performance in RoboTAP, while PO remains unchanged, suggesting that memory capacity alone is less critical than providing sufficiently long training sequences. Finally, we note that the full potential of memory is unlocked when it is extended at inference time, which we discuss in the next section.
TABLE VI: Training video length and memory size. We compare different training video lengths $T$ and memory sizes $L$, without applying inference-time memory extension (i.e. $L=L_{i}$). All metrics are higher-is-better.
$T$ $L$ DAVIS RoboTAP PO
@@
(i) 24 18 77.5 90.2 76.0 89.0 46.8
(ii) 36 18 78.3 90.7 78.2 91.2 51.3
(iii) 48 18 79.0 90.7 79.1 92.2 51.5
(iv) 48 24 79.1 91.5 78.9 91.5 51.5
@@ 4.5.6 Training Video Frame Sampling
While we sample $T$ frames from the 120-frame videos in the dataset, the way these frames are selected can be important. For example, if the object or camera motion is slow, sampling only sequential frames (stride 1) may provide little useful signal for learning motion. To explore the optimal frame sampling strategy, we experiment with two sampling methods and different ratios, as shown in Table VII.
@@
4.5.7 Memory Update Strategies
While Track-On2 adopts a simple FIFO-style memory update during training, memory behavior at inference time can be more flexible, especially for long videos. The memory update strategy during training is constrained by the limited temporal span of the training clips, which are restricted to at most 48 frames due to GPU memory constraints. As a result, the memory design must remain simple and well-defined within this regime while still enabling reliable temporal learning. Since we cannot explicitly train on very long-term dependencies under these constraints, we adopt a straightforward FIFO update that refreshes the memory at every frame ($k_{1}=1$).
TABLE VII: Training video frame sampling. Comparison of uniform and random strategies with different ratios $r$, where $r$ controls the temporal crop size before selecting $T$ training frames. All metrics are higher-is-better.
@@ 4.5.7 Memory Update Strategies
However, inference on substantially longer videos allows us to explore alternative update strategies without retraining, provided the temporal ordering of memory entries is preserved. We evaluate different memory update mechanisms in Table VIII on DAVIS, the longer Dynamic Replica (DR), and the very long Point Odyssey (PO). We divide the fixed total memory length $L$ into two partitions, $L_{1}$ and $L_{2}$, updated every $k_{1}$ and $k_{2}$ frames respectively ($k_{2}>k_{1}$). Here, the $L_{2}$ partition acts as a long-term memory component, extending the temporal receptive field to approximately $L_{1}\cdot k_{1}+L_{2}\cdot k_{2}$ frames.
TABLE VIII: Memory update mechanisms. Comparison of different memory update strategies. The total memory length $L$ is fixed and split into short-term ($L_{1}$) and long-term ($L_{2}$) components, updated every $k_{1}$ and $k_{2}$ frames, respectively. Results are reported on DAVIS, Dynamic Replica (DR), and Point Odyssey (PO). All metrics are higher-is-better.
Ratio Period DAVIS DR PO
$L_{1}$ $L_{2}$ $k_{1}$ $k_{2}$ AJ $\delta^{x}_{avg}$ OA $\delta^{x}_{avg}$ $\delta^{x}_{avg}$ Survival
(i) 1 - 1 - 67.0 79.9 92.0 74.5 45.1 57.7
(ii) 1 - 2 - 66.4 79.0 91.6 74.8 43.9 55.0
(iii) $1/2$ $1/2$ 1 2 66.5 79.4 91.8 74.5 44.4 58.4
(iv) $1/2$ $1/2$ 1 4 66.3 79.2 91.5 74.4 46.4 58.9
(v) $2/3$ $1/3$ 1 2 66.6 79.6 91.8 74.8 45.8 58.5
(vi) $2/3$ $1/3$ 1 4 66.3 79.2 91.6 74.9 48.5 59.6
The first two rows of Table VIII use a single partition ($L_{2}=0$). Our default setting (row i), where $L_{1}=L$ and $k_{1}=1$, performs best on DAVIS, where fine-grained temporal propagation is critical for short-term accuracy. Conversely, on longer sequences, incorporating a long-term component consistently improves performance. Experiment (vi), i.e., updating $2L/3$ of the memory every frame and the remaining $L/3$ every 4 frames, improves survival on Point Odyssey by 1.9 points and $\delta^{x}_{avg}$on Dynamic Replica by 0.4 points. Furthermore, increasing the long-term period from $k_{2}=2$ to $k_{2}=4$ yields consistent gains across different splits (comparing (iii) vs. (iv) and (v) vs. (vi)).
Overall, these results suggest that while dense, short-term memory updates are preferable for short videos, incorporating a sparse long-term memory component is beneficial for long-horizon tracking. This analysis highlights that memory in Track-On2 functions as a learned temporal mechanism whose effective behavior depends not only on its size but also on how it is updated over time.
Figure 16: Failure cases on uniform regions. On textureless or repetitive areas (highlighted regions), the spatial structure of uniformly initialized query points degrades over time due to ambiguous appearance-based matching. In contrast, points located on distinctive textured regions (e.g., the red discs) remain stable.
4.6 Failure Cases
Track-On2 primarily relies on appearance similarity for patch classification, while temporal information is incorporated implicitly through the memory mechanism. Unlike iterative update-based trackers, no explicit spatial location refinement or direct coordinate propagation is enforced across time steps. Although this design improves robustness to error accumulation and enables recovery from drift, it may introduce spatial instability under certain conditions.
In particular, predictions can exhibit jitter or drift in regions with uniform or repetitive texture, where appearance-based matching becomes inherently ambiguous. As illustrated in Fig. 16, we visualize uniformly initialized grid points on a RoboTAP sequence, sampled every 30 frames. While points located on distinctive textured regions (the red discs) remain stable, the spatial structure of points on textureless regions (highlighted in red rectangles) gradually degrades. This behavior reflects the difficulty of resolving correspondences in areas lacking discriminative visual cues. Consequently, Track-On2 is most reliable when tracking anchor keypoints with distinctive appearance, whereas uniformly distributed points over textureless regions are more prone to instability.
Another challenging scenario involves thin structures. Due to spatial resolution reduction in intermediate feature maps, fine geometric details may not be sufficiently preserved. As a result, the extracted query feature may not accurately represent the underlying pixel-level structure, leading to imprecise matching along thin surfaces.
@@ 5 Conclusion
Limitations and Future Work: While several design choices in Track-On2 naturally improve efficiency (e.g., strict frame-by-frame processing, a single expandable memory, and the absence of heavy 3D cost volumes), the model is not explicitly engineered for hardware-specific optimization. In particular, we rely on standard Transformer operations without tailoring low-level implementations for reduced FLOPs or memory footprint (e.g., attention kernel redesign, tensor operation fusion, or hardware-aware concatenation strategies), unlike recent object tracking works that optimize both architectural design and low-level operations for edge deployment [57, 58, 59]. As a result, although Track-On2 achieves competitive runtime on high-end GPUs (all benchmarks are conducted on NVIDIA A100), it has not yet been optimized for resource-constrained environments such as CPUs or embedded devices. Extending Track-On2 with hardware-aware optimization and system-level efficiency improvements is therefore a valuable direction for future work, particularly given the practical importance of real-time tracking in embedded and real-world applications.
Beyond efficiency, the failure cases discussed above highlight the limitations of relying predominantly on appearance-based matching. Explicitly modeling motion within the architecture is a promising direction for future work, as it could improve temporal coherence while reducing over-reliance on appearance. Future work also includes exploring more advanced memory update strategies beyond FIFO, such as adaptive memory writes conditioned on video length or query content, to better balance short-term and long-term information in very long sequences. More broadly, effective adaptation to real-world data remains an important challenge for point tracking models. Exploring real-world fine-tuning and improving pseudo-labeling strategies are key directions for enhancing robustness and generalization beyond synthetic training regimes.
@@ 6 Acknowledgements
This project is funded by the European Union (ERC, ENSURE, 101116486) with additional compute support from Leonardo Booster (EuroHPC Joint Undertaking, EHPC-AI-2024A05-028). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Weidi Xie would like to acknowledge the Scientific Research Innovation Capability Support Project for Young Faculty (ZYGXQNJSKYCXNLZCXM-I22).